what is the slope=3 and passes througth (5,-2).so what is the solutions?
Answer by scott8148(6628)

y = mx+b is the general slope-intercept; m = slope and b = y-intercept

for this case; -2 = 3(5)+b

-2 = 15+b ... subtracting 15 ... -17 = b

so the equation is y = 3x-17